Shift Summary - Sept 8, 1999 Day

Douglas A. Jensen (djensen@fnal.gov)
Wed, 8 Sep 1999 15:39:06 -0500 (CDT)

Shift Summary for Dat shift, 09/08/99

On Shift: Wah, D. Jensen

Hours of beam available at all intensities: 3:20 / 8:30
Hours of beam available at nominal intensity: 3:20 / 8:30
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 2:40

Accounting of lost time:

tevatron down due to MI and sextupole power supply problems.
========================================================================
Runs ended this shift: 14464 - no beam

Tapes brought to FCC this shift: NONE
(including tapes for online split)

========================================================================
Accesses during shift:

1. Move regenerator to way out. Unscrewed it. control cable unplugged
and program left running - to keep absorber moving, as per EB.
After a bit of a muddled start ( see below ) start taking lots
of Regenerator-OUT data.

========================================================================
Overview of shift:
waiting for beam to return. And waiting, and waiting, and......
Finally got some beam, running regenerator out data
Lost some time due to a false start - absorber not moving.
But recovered from that problem with the loss of about 30 min.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

07:00 shift crew arrives and discovers that tevatron
has been down since 23:11 due to MI problems.
estimate was 3 hours, start time of the 3 hrs
updating.


14465 08:45 Shift crew very optimistically begins a new run
as a contingency against the FANTASTICALLY unlikely
scenario that beam will actually return during
their shift. ( this remark copied exactly from
prev shift, except it is 5 hours later )
Regenerator WAY OUT. Goal, about 4 M Trig 6
= 3pi0. Est about 3 hrs needed. That is
assuming the beam is on. Seems like that might
be an irrational assumption.....

09:16 Est - 2 hours.
11:12 "No est"
11:45 MCR called to say they will try to establish beam to
us as soon as they are able to establish clearances
etc.
12:07 MCCR says they are now really going to send beam.
154 12:10 getting some beam
184 12:50 Realized, from DC plots, that absorber was not moving.
This is because epicure moves regenerator, and abs
follows. Discover there is a program ABSMOVE which
moves just the absorber. That seems to work.
Also, Elizabeth turns off reg HV at this time. No
use having it on. This also gets regenerator out of
veto. About 340K events to be ignored.

15:30 Hand over running experiment to the next shift
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
Put reg back in when appropriate.